OpenStudy (anonymous):

PLEASE HELP!! Somebody please explain to me how LCM(Lowest Common Multiple) works~~ Thanks!! :)

OpenStudy (anonymous):

sure. lets pick easy numbers and see how it is done. first of all math does not waste words so your least common multiple must be a) a multiple of both numbers b) the smallest one

OpenStudy (anonymous):

so if i take some nice small numbers like 6 and 8 i can just guess it is 24 because both 6 and 8 go in to 24 evenly, and 24 is the least number. multiples of 6 are 6, 12, 18, 24, 30, 36, .. multiples of 8 are 8, 16, 24, 32, 40 and you can see that 24 is in both lists so it is "common" and it is the smallest one

OpenStudy (anonymous):

What about 32 & 45? I know there are stupid ways of figuring this out, but it's not very efficient and I have a Math quiz tomorrow.

OpenStudy (anonymous):

ok now it gets a little more complicated. but not in this cases because \[32=2^5\] and \[45=3^2\times 5\] and you see they have no factors in common. in math we say they are "relatively prime"

OpenStudy (anonymous):

so your LCM is the product of the two numbers, namely \[32\times 45=1440\]

OpenStudy (anonymous):

12&14?

OpenStudy (anonymous):

all the work is in factoring the numbers to see which if any factors they have in common. lets try 12 and 14

OpenStudy (anonymous):

ok we factor 12 as \[12=2^2\times 3\] \[14=2\times 7\] so we are going to need \[2^2\] and \[3\] and \[7\] for our least common multiple we do not need \[2^3\] because if we have \[2^2\] we are good for both 17 and 14

OpenStudy (anonymous):

sorry for 12 and 14

OpenStudy (anonymous):

would help if i could multiply our least common multiple is \[2^2\times 3\times 7=4\times 3\times 7=12\times 7=84\]

OpenStudy (anonymous):

it is a multiple of 12 because \[12\times 7=84\] and it is a multiple of 14 because \[6\times 14=84\]

OpenStudy (anonymous):

Okay, yes, I understand that part. What confuses me is what numbers you should take out of the prime factors to figure out the LCM.

OpenStudy (anonymous):

ok all the work is in factoring the numbers. then you take each prime you see to the highest power you see it in any ONE number

OpenStudy (anonymous):

for example if i have \[2^2\times 3^4\times 7\] and \[2^3\times 3\times 7^2\] my LCM is \[2^3\times 3^4\times 7^2\]

OpenStudy (anonymous):

some times it is obvious. like if i had 30 and 50 i would say 150 right away

OpenStudy (anonymous):

OH! Yeah, now I get what prime factors you use to determine the LCM. OK. Thanks!

OpenStudy (anonymous):

good. each prime you see to the highest power you see it in any ONE factor. don't go adding up the exponents!

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

What if you have one with powers and another one without any powers? Just integers?

OpenStudy (anonymous):

if you have \[3\times 5^2\] and \[3\times 5\times 7\] you need \[3\times 5^2\times 7\]

OpenStudy (anonymous):

there isn't really "one without powers" if you just see a number the exponent is 1 \[3=3^1\]